OpenStudy (anonymous):

Using L'Hopital's rule,Evaluate the limit. lim as x approaches pi/4 9 tan u - 9 cot u divided by 4u-pi

OpenStudy (anonymous):

Wouldn't this be the limit as u approaches pi/4?

OpenStudy (anonymous):

still need this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you know what l'hopital's rule is? ^_^

OpenStudy (anonymous):

yes.. If the problem goes to 0/0 or infinity / infinity you take the derivative of both the numerator and denominator..

OpenStudy (anonymous):

yes. so we have: \[9 \lim_{u -> \pi/4}\frac{ \sec^2(u) + \csc^2(u) }{ 4 } = 9 *\frac{ \frac{ 1 }{ (\cos(\pi/4))^2 } +\frac{ 1 }{ (\sin(\pi/4))^2 } }{ 4 } = 9*\frac{ 2+2 }{ 4 } = 9\]

OpenStudy (anonymous):

can explain to me step by step on how you did this?

OpenStudy (anonymous):

can you help me with another problem?

OpenStudy (anonymous):

we started with \[9 \lim_{u ->\pi/4}\frac{ \tan(u) + \cot(u) }{ 4u - \pi }\]since this is 0/0 we take the derivative of them respectively [not to be confused with quotient rule] derivative of tanu = sec^2(u) derivative of cotu = -csc^2(u) derivative of denominator is 4 and sec^2(u) = (sec(u))^2 and sec(u) = 1/cos(u) csc(u) = 1/sin(u). i used a calculator for sin and cos of pi/4 [even though don't have to]

OpenStudy (anonymous):

i have to eat but if nobody helps by the time im done (like 10 minutes) i will

OpenStudy (anonymous):

okay .. Thanks..

OpenStudy (anonymous):

...

OpenStudy (anonymous):

I posted the other question. .

OpenStudy (anonymous):

limx→0 14csc^2(x)−14cot^2(x) using L'Hopital's rule..